OpenStudy (anonymous):

What is the fifth term of the sequence whose first term is 10 and whose common ratio is 1/2? A.160 B.50 C.12 D.5/8

OpenStudy (ash2326):

nth term of an arithmetic sequence is given as \[a_n=a_1r^{(n-1)}\] a_1= first term r= common ratio a_n= nth term here n=5 \(a_1=10\) and common ratio \(=\frac{1}{2}\), can you find the fifth term now?

Parth (parthkohli):

Think of it like this. First term - 10 Second - 10 * 1/2 Third - (10 * 1/2) * 1/2 Fourth - (10 * 1/2 * 1/2) * 1/2 Fifth - (10 * 1/2 * 1/2 * 1/2) * 1/2

Parth (parthkohli):

So first find 1/2 * 1/2 * 1/2 * 1/2. \( \color{Black}{\Rightarrow ({1 \over 2})^4 = {1 \over 2^4} = {1 \over 16}}\) Now find 10 * 1/16

OpenStudy (anonymous):

is it 5/8?

Parth (parthkohli):

Yep